an electron (me = 9.10939 × 10–28 g) moving at 3.90 × 106 m/s in an electron microscope. Answer in nm.
1
Expert's answer
2019-09-30T04:54:23-0400

λ = h / p = h / m x v=6.62 x 10 -34 / 9.10939 x 10-28 x 3.9x106=1.8633914e-13 m.

0.00018633914 nm